OpenStudy (anonymous):

Let f(x)=(x+3)/7. show that f^-1=7x-3. show work. would the answer just be the inverse steps?

OpenStudy (anonymous):

f(x)=y interchange x and y solve for y replace y with f^-1(x)

OpenStudy (freckles):

Solve y=(x+3)/7 for x

OpenStudy (freckles):

or you know nevermind we are trying to prove

OpenStudy (freckles):

check to see if this is true: \[f(f^{-1}(x))=f^{-1}(f(x))\]

OpenStudy (freckles):

=x

OpenStudy (freckles):

\[f(f^{-1}(x))=f^{-1}(f(x))=x\]

OpenStudy (anonymous):

sorry i should of put the question for the function f use composition of functions to show the f^-1 is as given. You may use composition of fuctions or you may find a formula for the inverse.

OpenStudy (freckles):

compositions is what i have above

OpenStudy (freckles):

f^-1 is 7x-3 so to find f(f^-1(x)) replace first f^-1 with 7x-3 f(7x-3) replace the x in f with 7x-3

OpenStudy (freckles):

and you get x back then you win! :)

OpenStudy (freckles):

\[f(f^{-1}(x))=f(7x-3)=\]you need to show this is x

OpenStudy (anonymous):

where does the 7x-3 go in the f^-1(x)?

OpenStudy (freckles):

f^{-1} is 7x-3 so i just replace f^(-1) with 7x-3

OpenStudy (freckles):

you need to plug 7x-3 into f now where x is

OpenStudy (freckles):